LControl

LANC-/5-Pin-Edit-Interface for Amiga Computers

Version 1.51 June 1995

by Juergen Frank und Michael Watzl

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993-95 Juergen Frank und Michael Watzl

1 What is LControl

LControl is a program to control video cameras via a special control hardware. The cameras must have a LANC- (Sony compatible) or 5-Pin-Edit-interface (Panasonic compatible).

For cameras with a LANC-interface there are hardwares for the (internal) parallel and the (internal) serial port of the Amgia. The solution for the serial port needs much less CPU power. The parallel port hardware offers a two camera option – the software can swap between the two connected cameras.

For cameras with a 5-Pin-Edit-interface (Panasonic/Blaupunkt) there is a parallelport hardware, too. (LControl becomes PControl).

You will get the schematics for the hardware, the construction kit or the complete hardware^{1} after the registration. See Appendix B [Registering], page 9.

LControl is a commodity and it has an ARexx port. Therefore LControl can work in the background and your camera can be "programmed" via ARexx.

See also Appendix C [IR-Master-Support], page 10.

2 Installation

Requirements for LControl/SControl/PControl:

• Amiga computer

• Amiga-OS 2.04 or better

• a video camera with - LANC-interface (for LControl/SControl)

- 5-Pin-Edit-interface (for PControl)

Start the installation process by doubleclicking the ‘LControl english’ icon. The installation script^{2} will run automatically.

1 For 5-Pin-Edit-cameras we only offer the schematics 2 Installer (c) 1991-1993 Commodore-Amiga, Inc. All Rights reserved. Reproduced and dis-

tributed under licence from Commodore. Installer Software is provided "as-is" and subject to

3 How To Use

3.1 Menus

3.1.1 Project

About... Gives some information about the authors and the program. Hide Closes the LControl window. However, LControl stays active. Fix position

Snapshots the current position. This piece of information is saved in the tooltypes. See Section 3.4 [Tooltypes], page 6.

Quit Quits LControl.

3.1.2 Special

Reset counter

Resets the counter of the camera.

Zoom in Sends "Zoom In" to the camera Zoom out Sends "Zoom Out" to the camera Record Sends "Record" to the camera. Eject Sends "Eject" to the camera. Gotomode In the submenu you can select one of these commands.

0 Play 1 Stop 2 Pause 3 Record The selected command is sent when a successful "goto" was made.

Goto You may enter a counter. Then the tape is winded to this position an the command

selected with gotomode is sent. change; no warranties are made. All use is at your own risk. No liability or responsibility is assumed.

3.1.3 Special-II

Note: All the commands in this menu refer directly to standard camera commands. Please look up their meaning in your camera’s user manual.

3.2 Gadgets

Swap Swaps to the other camera^{3}. Play Sends a "Play"-command to the camera. Stop Sends a "Stop"-command to the camera. Pause Sends a "Pause"-command to the camera. FREW Sends a "fast rewind" command to the camera. FFOR Sends a "fast forward" to the camera. PREW Sends a "picture rewind" to the camera. PFOR Sends a "picture forward" to the camera. −− Sends a "slow motion rewind" to the camera. − Sends a "single pic back" to the camera. + Sends a "single pic forward" to the camera. ++ Sends a "slow motion forward" to the camera.

3.3 ARexx Port

The name of the ARexx port of

LControl (LANC/parallel) is LControl_rexx SControl (LANC/serial) is SControl_rexx PControl (5-Pin-Edit/parallel) is PControl_rexx

So far, the following commands are supported: 3 LControl (parallel) only ‘LC_QUIT’ Quits LControl. ‘LC_SHOW’ Opens the LControl window. ‘LC_HIDE’ Closese the LControl window.

‘LC_ENABLE’

Opens the LControl window and opens the ARexx port (if closed).

‘LC_DISABLE’

Closes the LControl window and the ARexx port.

‘LC_PLAY’
‘LC_STOP’
‘LC_PAUSE’
‘LC_FREW’
‘LC_FFOR’
‘LC_PREW’
‘LC_PFOR’
‘LC_PLUS’
‘LC_MINUS’
‘LC_SLOW’
‘LC_RSLOW’
‘LC_GETCOUNTER’

The current counter will be returned in RESULT.

‘LC_GETSTATUS’

Returns the current status in RESULT:

NO CAMERA

If no camera/hardware is connected.

CAMERA0: <action>

When camera 0 is the active one.

CAMERA1: <action>

When camera 1 is the active one.

‘LC_RESET’

Resets the counter of the camera.

‘LC_SWAP’ Swaps to the other camera^{4}.

‘LC_ZOOMIN’

Zooms in one step.

4 Only for LControl/parallel

‘LC_ZOOMOUT’

Zooms out one step.

‘LC_RECORD’

Sends Record to the camera.

‘LC_EJECT’

Sends Eject to the camera.

‘LC_GOTOMODE’

Expects a number (0,1,2 or 3) as parameter. See Section 3.1.3 [Menus], page 3.

‘LC_GOTO’ Expects a counter as parameter. See Section 3.1.3 [Menus], page 3.

‘LC_INSAUDIO’
‘LC_INSVIDEO’
‘LC_ASSEMBLE’
‘LC_INSA_LON’
‘LC_INSA_LOFF’
‘LC_INSA_RON’
‘LC_INSA_ROFF’
‘LC_INSV_ON’
‘LC_INSV_OFF’
‘LC_ASSEMBLEON’
‘LC_ASSEMBLEOFF’
‘LC_RAWCOMMAND’

This command directly sends data to the camera.

WARNING: Use this command only when you really know what you are doing!

Expects hexadezimal bytes as parameters. Example:

               LC_Rawcommand ff 10 aa

This will send 255, 16 and 170 to your camera.

     Basic VTR-commands

hex                 command
00                  CH-1/1
40                  CH-2/2
20                  CH-3/3
60                  CH-4/4
10                  CH-5/5
50                  CH-6/6
30                  CH-7/7
70                  CH-8/8
08                  CH-9/9
48                  CH-10/10
28                  CH-11
68                  CH-12/CH/ENTER/#
18                  CH-13/1-
58                  CH-14/2-
38                  CH-15
78                  CH-16
04                  CH-HIGH
44                  CH-LOW
14                  X3 OR 2X
54                  POWER ON/OFF
22                  SLOW 2
62                  SLOW 3
32                  HIGH DOUBLE SPEED
2A                  ANT-SW
5A                  ANT-VTR
3A                  POWER ON
7A                  POWER OFF
06                  REVERSE
46                  FORWARD
09                  INDEX WRITE
49                  INDEX ERASE
65                  INDEX

3.4 Tooltypes

The following tooltypes are recognized:

CX POPUP= "YES" or "NO"

YES Pop window up on startup

NO Start hidden CX POPKEY=

Hotkey for LControl e.g.: lshift f10 ↦ left shift-key + F10

Default: lcommand esc.

LEFT=xxx Leftedge of the window

TOP=yyy Topedge of the window PUBSCREEN=name

If specified, LControl tries to open its gui on the publicscreen named ‘name’.

GOTOMODE= Expects a decimal number (0,1,2 or 3) as parameter. See Section 3.1.3 [Menus], page 3, gotomode. You can specify the default command which should be sent when positioning with the goto-function.

LIBNAME= Here you can specify the library name which should be used.

LANC For the parallelport hardware for LANC-cameras (LControl).

SLANC For the serialport hardware for LANC-cameras (SControl).

EDIT For the parallelport hardaware for 5-Pin-Edit-cameras (PControl).

DEBUG Just fakes a camera (DControl).

FAST AT= If the distance between the current counter and the target is more than <fast at> seconds for a goto, LControl should use FFOR/FFREW for positioning.

PLAY AT= If the distance between the current counter and the target is less than <play at> seconds for a goto, LControl should use only PLAY for positioning.

SAVEPLAY= If the distance betwenn the current counter and the target is even less than <saveplay> seconds for a goto, LControl should rewind first.

FFOR OFFSET= FREW OFFSET= PFOR OFFSET= PREW OFFSET=

If the camera is in ths mode xxxx and a STOP is sent, then the xxxx OFFSET is the differece between the counter where the STOP was sent and the counter where the camera really stops.

Appendix A Support & Bugreports

A.1 Support

You may simply send a disk and enough stamps to one of the following addresses to get the newest software version of LControl:

Michael Watzl               or      Jürgen Frank
Konradstr. 11                       Wittelsbacherweg 7
86609 Donauwörth                    86609 Donauwörth
Tel.: +49 906 5834                  Tel.: +49 906 1057
Germany                             Germany

A.2 Bugreports

Bugreports or suggestions are welcome and should be sent to one of the addresses above mentioned.

Please include a brief description of your computer and its configuration — this will help us to reproduce the bugs.

Bugreports and suggestions may also be sent to Michael Watzl via e-mail (indy@abs.gun.de).

Appendix B Registering (1.51)

   NAME:       ___________________________________________
   STREET:     ___________________________________________
   ZIP-CODE:   ______CITY:     ________________________
   COUNTRY:    ____________________TEL: _________________
I want to become registerd user of LControl/SControl or
PControl, therefore I
o include a cheque
o include a money order
o include cash
Please send me:
o newest version + schematics           o newest version + schematics
  (par/ser for LANC-cameras)              (par for 5-Pin-Edit hardware)
  for $12 / DM 20                         for $12 / DM 20
  newest version + construction kit
o serial    LANC/Control-L hardware
o parallel  LANC/Control-L hardware
  for $18 / DM 30
  newest version + complete hardware
o for serial port LANC/Control-L
o for parallel port LANC/Control-L
  for $30 / DM 40
-----  Optional [for construction kit LANC/Control-L] -----------------------
o two-camera-set [allows connection of two cameras to LControl hardware]
  plus $3 / DM 5
-----  Optional [for complete hardware, parallel LANC/Control-L] ------------
o two-camera-set [allows connection of two cameras to LControl hardware]
  plus $6 / DM 10
-----------------------------------------------------------------------------
Together  __________$ / Mark
_______________________________________________
 locality, date                            sign

Appendix C IR-Master-Support

Taken from "IR-Master.guide": "The IR-Master package consists of two software parts and a control hardware which is plugged to the joystickport of your Amiga. With this equipment almost any device shipped with an IR remote can be controlled by your Amiga. At first you edit a virtual IR remote with the IR-Editor. Then this project is saved to disk an launched with the IR-Runner. The number of active IR-Runners is not limited. You can use the IR-Runner either interactively or control it via ARexx from other applications or even programm it using ARexx scripts. There are also some optional (send-only) hardwares: One is connected to the (internal) parallel port of the Amiga and carries up to eight IR senders. This eight channels can be selectively used by the IR-Master. Another optional hardware is designed for the audio-port. Here, the sending of IR commands uses DMA^{1} with the advantage not to block the system for a moment. With IR-Master 3.20 a new hardware for the (internal) serial port of the Amiga is introduced. LControl and other applications to control video cameras and players are supported directly since version 2.5 of the IR-Master."

1 direct memory access – does not stress the CPU

Appendix D History

V1.00
first Release
V1.01
- arexx-command: GETSTATUS added
- fixed some minor bugs
V1.02
- Support for 2 cameras added [Switch]
- minor bugfixes
V1.05
- Zoom in/Zoom out added
- Record added
V1.1
- Goto added
- Better SWAP implementation
V1.2
[never released]
V1.3
- Now LControl uses a library (LANC.library)
- goto added (automatic positioning)
- many other minor bugfixes an features
V1.4
- localized
V1.50
- pos-routine improved
- camera parameters can be modified to
  optimize pos-routine
- multiple configurations supported via project icons

V1.51
- calling pause in status record did send
  rubbish to the player

Index 12

Index

-

-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

– . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

+

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ++ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

5

5-Pin-Edit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

A

About... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Amiga-OS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 ARexx Port. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Authors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

B

Blaupunkt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

C

Control-L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 CX POPKEY= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 CX POPUP= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

D

DEBUG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 DEBUG.library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

E

E-Mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 EDIT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 EDIT.library. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Eject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

F

FAST AT= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 FFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 FFOR OFFSET= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Fix position. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 FREW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 FREW OFFSET= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

G

Goto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Gotomode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 GOTOMODE=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

H

Hardware . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Hide . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 How to use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

I

Infra Red . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Installer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 IR-Master-Support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

L

LANC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 7 LANC.library . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 LC ASSEMBLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC ASSEMBLEOFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC ASSEMBLEON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC DISABLE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC EJECT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC ENABLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC FFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC FREW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC GETCOUNTER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC GETSTATUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC GOTO. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC GOTOMODE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC HIDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC INSA LOFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSA LON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Index 13

LC INSA ROFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSA RON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSAUDIO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSV OFF. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSV ON . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC INSVIDEO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC MINUS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PAUSE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PLAY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PLUS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC PREW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC QUIT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC RAWCOMMAND. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC RECORD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LC RESET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC RSLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SHOW. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SLOW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC STOP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC SWAP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC ZOOMIN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 LC ZOOMOUT. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 LEFT=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 LIBNAME= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

M

Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

Panasonic. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Parallel port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Pause. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PFOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PFOR OFFSET= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Play . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PLAY AT=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 PREW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 PREW OFFSET= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Prices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 PUBSCREEN=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Q

Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

R

Record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Registering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Reset counter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

S

SAVEPLAY=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 SLANC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 SLANC.library. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Sony. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Special . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Special-II . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Swap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 TOP= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

U

Update . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

V

Videocamera. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

W

What is LControl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Z

Zoom in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Zoom out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Table of Contents

1 What is LControl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 3 How To Use. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3.1 Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.1.1 Project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.1.2 Special. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.1.3 Special-II . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.2 Gadgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.3 ARexx Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.4 Tooltypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Appendix A Support & Bugreports. . . . . . . . . . . . . . . . . . . 8

A.1 Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 A.2 Bugreports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

Appendix B Registering (1.51) . . . . . . . . . . . . . . . . . . . . . . . . 9 Appendix C IR-Master-Support . . . . . . . . . . . . . . . . . . . . . 10 Appendix D History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12